OpenStudy (swissgirl):

Prove: The set \( T=\{ n \in \mathbb{Z} : n=2(mod6) \} \) is denumerable

OpenStudy (swissgirl):

Is the following proof correct?

OpenStudy (swissgirl):

Define a function F on the integers by setting F(z)= 6z+20. F is one to one because if F(u)=F(v) then 6u+20 = 6v+20 so u=v. Every element t of T has the form 6k+2 for some integer k and t=F(k-3) so F maps onto T. Therefore T is equivalent to \( |mathbb{Z} \) which is denumerable

OpenStudy (swissgirl):

@KingGeorge can you check this when you have the time. Now that i reread it i think its correct but i just want ur opinion on this

OpenStudy (kinggeorge):

I believe your proof is correct.

OpenStudy (swissgirl):

ya but y cldnt we just have \(\huge f(z) = \frac {z-2} {6} \) ?

OpenStudy (swissgirl):

I am just wondering if we can choose nething?

OpenStudy (kinggeorge):

You can't quite choose anything. The example you just posted is one you probably should avoid since it doesn't output an integer for all \(z\in\mathbb{Z}\).

OpenStudy (swissgirl):

For example like what?

OpenStudy (kinggeorge):

Let \(z=3\). Then \(f(3)=\frac{3-2}{6}=\frac{1}{6}\)

OpenStudy (swissgirl):

ohhhh that is cool

OpenStudy (kinggeorge):

Unless of course you defined the function as follows \[f\;:T\longrightarrow\mathbb{Z}\]\[\quad \;\;\,z\longmapsto\frac{z-2}{6}\]This function would be fine to use.

OpenStudy (swissgirl):

ohhhhh okkkkk i seeeeee

OpenStudy (swissgirl):

Thanks you are awesome

OpenStudy (kinggeorge):

You're welcome.